05 February 2008

RHIC pp Set-Up:  Experiments got collisions last night.  STAR saw a global timing shift of 17 ns with lots of background.  PHENIX saw a 1ns shift, and 15kHz of Beam-Beam counters coincidence counts.  STAR saw 37kHz with spikes 3 times per minutes where the instantaneous rate from the Blue beam jump up by a factor of 50.  This correlates with a Beam Loss of 200% per hour.  Yellow beam looks great.  Blue beam is very touchy, shows a 1 minute oscillation in Wall Current Monitor readings.  Haixin made a measurement of the Blue beam emittance and saw 27pi mm-mr.  This is quite large, but agrees with emittance values derived from the STAR event rate.  Lowering the tunes by 0.02 improves the Blue orbit a lot (Todd) but we don't know what effect this might have on polarization.  We need to put the Blue Vertical Damper back in.

Last night, just as RHIC was ready to give beam to the experiments, the PHENIX Pad Chamber HV power supplies tripped off.  They will need an access to fix the problem. 

There was a quench link interlock at about 9:00AM this morning.  This appears to be due to a power supply failure.  As experts go in to diagnose, PHENIX will take the opportunity to take an access to work on their Pad Chamber HV power supplies. 

Today we will explore tunes around 0.91. 

Repair Yellow BBQ. Do feedback ramp.

Ramp tuning for transmission, store lifetime, polarization.

Polarization measurements.

Rotator ramp development (Waldo) unless we change the Blue ramp in which case the rotator ramp would be postponed until after the Blue ramp is completed and retuned. 

 


Scheduling Physicist: Mike Sivertz Last updated:
Email: sivertz@bnl.gov
Office Phone: 631-344-6102
Digital Pager: 7528 or 4300
Home Phone: 631-878-9020